-27=8x-5x
and
-37=3x+11-7x

Answer by checkley77(12844)

-27=8x-5x
3x=-27
x=-27/3
x=-9 answer.
-------------------------
-37=3x+11-7x
3x-7x=-37-11
-4x=-48
x=-48/-4
x=12 answer.